Common Ion Effect: Calculate Solubility

1) The maximum amount of silver chloride that will dissolve in a 0.204 M silver acetate solution is _______M.

2) The molar solubility of barium sulfate in a 0.164 M sodium sulfate solution is ________ M.
Answer #1

1)

Ksp ( AgCl) = [Ag+] * [Cl-]

suppose the solubility of silver chloride = S M

[Ag+] = S + 0.204

[Cl-] = S

1.6*10^-10 = [ S + 0.204] * S

since S + 0.204 is very less compared to 0.204

so S + 0.204 ~ 0.204

1.6*10^-10 = 0.204 * S
S = 1.6*10^-10 / 0.204

S = 7.843*10^-10 M

so The maximum amount of silver chloride that will dissolve in a 0.204 M silver acetate solution is 7.843*10^-10 M